OpenStudy (minecraftqueen):

Write the following expression using a single exponent. (25)2 × 27 249 23 214 217

OpenStudy (minecraftqueen):

(2^5)^2(2^7) =(32^2)(2^7) =1024(2^7) =(1024)(128) =131072

OpenStudy (minecraftqueen):

only problem is thats not up there so im lost

OpenStudy (minecraftqueen):

@Hero

OpenStudy (anonymous):

\[25^2 * 27=16875 \]

OpenStudy (minecraftqueen):

.... okie zz how dose that help

OpenStudy (superdavesuper):

here: http://mathforum.org/library/drmath/view/55602.html u got the right ans already :)

OpenStudy (minecraftqueen):

nvm test timed out agen -_-

OpenStudy (superdavesuper):

@robtobey the Q is actually (2^5)^2.2^7 ;)

OpenStudy (minecraftqueen):

thx any way tho

OpenStudy (anonymous):

\[131072=2^{17} \]